Need help solving:
-3x +3/4=1/8
Thanks

Answer by fractalier(6550)

From -3x + 3/4 = 1/8
we need to to subtract 3/4 from both sides
-3x = 1/8 - 3/4
-3x = -5/8
then divide both sides by -3
x = 5/24